# Adapting Newton’s Method to Neural Networks through a Summary of Higher-Order Derivatives

### v3

@@ 6 Discussion
Method 1 does not come with a precise convergence rate. The rate proposed in Appendix F (Theorem 1) gives only a heuristic. Given the convergence rates of Newton’s method and Cauchy’s steepest descent, we can expect to find some in-between convergence rates. Since Cauchy’s steepest method is vulnerable to a highly anisotropic Hessian, it would be valuable to know how much this weakness is overcome by our method.
Practicality.
Despite the interesting properties of Method 1 (scalability, invariance by reparameterization, evaluation of long-range interactions between layers), we have proposed nothing more than a proof-of-concept. This method remains subject to instabilities during training, which is expected for a second-order method, but not acceptable for the end user. Therefore, some additional tricks should be added to improve the stability of the training, which is a common practice, but usually comes with additional hyperparameters to tune. Also, improvements are needed to reduce the duration of each epoch, which is longer than with K-FAC.
Appendix A Extensions of Pearlmutter’s trick
In this appendix, we show how to use the trick of Pearlmutter (1994) to compute the terms of the Taylor expansion of $\mathcal{L}$ and the tensor $\mathbf{D}^{d}_{\bm{\theta}}(\mathbf{u})$ defined in Eqn. 6.
@@ Appendix A Extensions of Pearlmutter’s trick
$\mathcal{L}(\bm{\theta}+\mathbf{u})=\mathcal{L}(\bm{\theta})+\sum_{d=1}^{D}\frac{1}{d!}\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})[\mathbf{u},\cdots,\mathbf{u}]+o(\|\mathbf{u}\|^{D}).$ (11)
@@
$\tilde{\mathbf{D}}_{\bm{\theta}}^{d}(\mathbf{u}):=\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})[\mathbf{u},\cdots,\mathbf{u}]\in\mathbb{R}.$ (12)
@@
$\tilde{\mathbf{D}}_{\bm{\theta}}^{d-1}(\mathbf{u})=\frac{\mathrm{d}^{d-1}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d-1}}(\bm{\theta})[\mathbf{u},\cdots,\mathbf{u}]\in\mathbb{R}.$ (13)
@@
$\tilde{\mathbf{D}}_{\bm{\theta}}^{d}(\mathbf{u})=\left(\frac{\mathrm{d}\tilde{\mathbf{D}}_{\bm{\theta}}^{d-1}(\mathbf{u})}{\mathrm{d}\bm{\theta}}\right)^{T}\mathbf{u}.$ (14)
Therefore, at each step $d$, we only have to compute the gradient of a scalar $\tilde{\mathbf{D}}_{\bm{\theta}}^{d}(\mathbf{u})$ according to $\bm{\theta}\in\mathbb{R}^{P}$, and compute a dot product in the space $\mathbb{R}^{P}$. So, computing $\tilde{\mathbf{D}}_{\bm{\theta}}^{d}(\mathbf{u})$ has a complexity proportional to $d\times P$, and does not require the computation of the full tensor $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}(\bm{\theta})\in\mathbb{R}^{P^{d}}$.
Fast computation of $\mathbf{D}^{d}_{\bm{\theta}}(\mathbf{u})$.
We assume that the parameter $\bm{\theta}$ is represented by a sequence of vectors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$: each coordinate $\theta_{i}$ belongs to exactly one of the $\mathbf{T}^{s}$. In the same way, given a direction $\mathbf{u}\in\mathbb{R}^{P}$ in the space of the parameters, $\mathbf{u}$ can be represented by a sequence of vectors $(\mathbf{U}^{1},\cdots,\mathbf{U}^{S})$.
We want to compute the tensor $\mathbf{D}^{d}_{\bm{\theta}}(\mathbf{u})\in\mathbb{R}^{S^{d}}$, whose coefficients are:
$(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{s_{1},\cdots,s_{d}}$ $=\frac{\partial^{d}\mathcal{L}}{\partial\mathbf{T}^{s_{1}}\cdots\partial\mathbf{T}^{s_{d}}}(\bm{\theta})[\mathbf{U}^{s_{1}},\cdots,\mathbf{U}^{s_{d}}],$ (15)
@@
Let us assume that $\mathbf{D}^{d-1}_{\bm{\theta}}(\mathbf{u})$ is available. We can compute the coefficients of $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ as follows:
@@
$(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{s_{1},\cdots,s_{d}}=\left(\frac{\partial(\mathbf{D}^{d-1}_{\bm{\theta}}(\mathbf{u}))_{s_{1},\cdots,s_{d-1}}}{\partial\mathbf{T}^{s_{d}}}\right)^{T}\mathbf{U}^{s_{d}}$ (16)
@@
That way, the tensor $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ can be computed without having to compute the full derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\bm{\theta}^{d}}$. In fact, we do not need to store objects of size greater than $S^{d-1}\times P$: the last operation requires storing $\mathbf{D}^{d-1}_{\bm{\theta}}(\mathbf{u})$, which is of size $S^{d-1}$, and the gradient of each of its elements with respect to $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$, which is of size $P$.
Python implementation.
We provide here an example of implementation of this extension of Pearlmutter’s trick. First, we define in Alg. 2 the function dercon, standing for “derivation+contraction”. Second, we provide in Alg. 3 the function diff, which outputs a list of tensors $(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{d\in\{0,\cdots D\}}$.
We recall that $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})\in\mathbb{R}^{S^{d}}$, where $s$ is the number of groups of tensors in $\bm{\theta}$, $\mathbf{u}\in\mathbb{R}^{P}$ is a direction in the parameter space $\mathbb{R}^{P}$, that can be represented as a tuple of groups of tensors, with the exact same structure as $\bm{\theta}$.
Also, to keep this implementation efficient for $d\geq 2$, we reduce the computational and storage cost of the tensors $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$ by using the symmetry structure of $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})$. More specifically, for any permutation $\sigma:\{1,\cdots,d\}\rightarrow\{1,\cdots,d\}$:
$\forall(i_{1},\cdots,i_{d}),\quad(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{i_{1},\cdots,i_{d}}=(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{i_{\sigma(1)},\cdots,i_{\sigma(d)}}.$ (17)
So, instead of computing and storing the full tensor $\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u})\in\in\mathbb{R}^{S^{d}}$, we just have to store the coefficients $(\mathbf{D}_{\bm{\theta}}^{d}(\mathbf{u}))_{i_{1},\cdots,i_{d}}$, with $i_{1}\leq\cdots\leq i_{d}$. So, we compute and store only $\frac{(S+d-1)!}{d!(S-1)!}$ coefficients instead of $S^{d}$. For example, with $S=10$ groups of parameters and $d=3$, we store only $220$ coefficients instead of $S^{d}=1000$. We typically use this trick in Alg. 3.
Algorithm 2 Implementation of the dercon function, which performs a derivation+contraction operation. This operation allows us to gather higher-order information without having to store large tensors.
import torch
from . import select_params, dot
Existingfunctions:
*select_params(x,s0):
␣␣␣returnsthegroupsofparamsofxwithindicess>=s0
*dot(x1,x2):
␣␣␣x1andx2aretuplesoftuplesoftensorswiththesamestructure
␣␣␣␣␣a)computethedot-productbetweeneachpairoftensors(t1,t2)
␣␣␣␣␣b)sumtheseinsideenchgroupoftensors
␣␣␣␣␣c)returntheresult(whichisatorch.tensorcontainingas
␣␣␣␣␣manycoeffsastherearegroupsofparamsinx1andx2
def dercon(f, theta, u, s0):
# Derivation + contraction
# Derivation
theta_s0 = select_params(theta, s0)
deriv = torch.autograd.grad(f, theta_s0, create_graph = True)
# Contraction
u_s0 = select_params(u, s0)
result = dot(deriv, u_s0)
return result
Algorithm 3 Implementation of the diff function, computing $\mathbf{D}^{d}$ iteratively at several orders. We denote by L the loss function, and by (x, y) a data point.
import torch
from itertools import combinations_with_replacement
from . import dercon
def diff(L, x, y, order):
lstD = [None]*(order+1) # initialize the object to output
# Forward pass
loss = L(x, y)
# Store the loss (without keeping the computational graph)
lstD[0] = loss.detach()
# Compute the derivative of the loss and contract it
# in the directions given by the groups of params of u.
deriv = {tuple(): dercon(loss, u, 0)}
# Store the result (without keeping the computational graph)
lstD[1] = {k: v.detach() for k, v in deriv.items()}
# Computations with d >= 2
for d in range(2, order+1):
new_deriv = {}
# We only compute the derivatives of certain coefficients
[…]
# it is not necessary to compute all of them.
# Create the list of indices of the relevant coefficients
lst_idx = [tuple(sorted(idx)) for idx in \
combinations_with_replacement(range(S), d-1)]
# Compute these coefficients
for idx in lst_idx:
[…]
init, last = idx[:-1], idx[-1]
imax = last if len(init)==0 else last-init[-1]
# To access that coeff, we use deriv[init][imax]
new_deriv[idx] = dercon(deriv[init][imax], u, last)
# Store the result (without keeping the comp. graph)
lstD[d] = {k: v.detach() for k, v in new_deriv.items()}
# Prepare the next iteration
deriv = new_deriv
return lstD
Appendix B Derivation of the second-order method
We consider an update of $\bm{\theta}$ with one learning rate $\eta_{s}$ for each subset $\mathcal{I}_{s}$ of parameters. Let $\mathbf{I}_{S:P}\in\mathbb{R}^{S\times P}$ be the partition matrix, verifying $(\mathbf{I}_{S:P})_{sp}=1$ if $p\in\mathcal{I}_{s}$ and 00 otherwise, and $\mathbf{I}_{P:S}:=\mathbf{I}_{S:P}^{T}$. We consider an update based on a given direction $\mathbf{u}_{t}$ and we define $\mathbf{U}_{t}:=\mathrm{Diag}(\mathbf{u}_{t})$:
$\bm{\theta}_{t+1}=\bm{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta},$ (18)
where $\bm{\eta}=(\eta_{1},\cdots,\eta_{S})\in\mathbb{R}^{S}$.
@@ Appendix B Derivation of the second-order method
$\mathcal{L}(\bm{\theta}_{t+1})$ $=\mathcal{L}(\bm{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta})$
$=\mathcal{L}(\bm{\theta}_{t})-\bm{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta}_{t})$
$\hphantom{=}\;+\frac{1}{2}\bm{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\bm{\theta}^{2}}(\bm{\theta}_{t})\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}+o(\|\bm{\eta}\|^{2})$
$=\mathcal{L}(\bm{\theta}_{t})-\bm{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{g}_{t}$
$\hphantom{=}\;+\frac{1}{2}\bm{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{H}_{t}\mathbf{U}_{t}\mathbf{I}_{P:S}\bm{\eta}+o(\|\bm{\eta}\|^{2})$
$=\mathcal{L}(\bm{\theta}_{t})-\bm{\eta}^{T}\bar{\mathbf{g}}_{t}+\frac{1}{2}\bm{\eta}^{T}\bar{\mathbf{H}}_{t}\bm{\eta}+o(\|\bm{\eta}\|^{2}),$
@@
Now, we omit the $o(\|\bm{\eta}\|^{2})$ term and we want to minimize according to $\bm{\eta}$ the variation of the loss:
$\mathcal{L}(\bm{\theta}_{t+1})-\mathcal{L}(\bm{\theta}_{t})\approx\bm{\Delta}_{2}(\bm{\eta}):=\mathcal{L}(\bm{\theta}_{t})-\bm{\eta}^{T}\bar{\mathbf{g}}_{t}+\frac{1}{2}\bm{\eta}^{T}\bar{\mathbf{H}}_{t}\bm{\eta}.$ (20)
We have: $\frac{\mathrm{d}\bm{\Delta}_{2}}{\mathrm{d}\bm{\eta}}=-\bar{\mathbf{g}}_{t}+\bar{\mathbf{H}}_{t}\bm{\eta}$, which is zero if, and only if: $\bar{\mathbf{g}}_{t}=\bar{\mathbf{H}}_{t}\bm{\eta}$. If this linear system can be inverted, one can choose:
$\bm{\eta}=\bm{\eta}_{t}^{*}:=\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}.$ (21)
@@
This method can also be derived by taking the point of view of multivariate optimization. Within our setup, $\bm{\theta}$ is considered as a tuple of tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. Thus, we want to minimize the loss $\mathcal{L}$ with respect to $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. By abuse of notation, we will write:
$\mathcal{L}(\bm{\theta})=\mathcal{L}(\mathbf{T}^{1},\cdots,\mathbf{T}^{S}).$ (22)
@@
$\approx-\bm{\eta}^{T}\bar{\mathbf{g}}+\frac{1}{2}\bm{\eta}^{T}\bar{\mathbf{H}}\bm{\eta},$ (26)
where $\bm{\eta}=(\eta_{1},\cdots,\eta_{S})\in\mathbb{R}^{S}$, $\bar{\mathbf{g}}\in\mathbb{R}^{S}$ is the gradient of $f$ and $\bar{\mathbf{H}}\in\mathbb{R}^{S\times S}$ is the Hessian of $f$:
@@
Finally, one can minimize the order-2 approximation of $f$ (Eqn. 26) with respect to $\bm{\eta}$, with any numerical or analytical technique. If done analytically, we roll back to Eqn. 21.
@@
$g(\mathbf{v})=\mathcal{L}(\bm{\theta}-\mathbf{v})-\mathcal{L}(\bm{\theta}),$ (29)
@@
$g(\mathbf{v})\approx\tilde{g}(\mathbf{v}):=-\mathbf{v}^{T}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta})+\frac{1}{2}\mathbf{v}^{T}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\bm{\theta}^{2}}(\bm{\theta})\mathbf{v}.$
@@
$\mathbf{v}^{*}=\operatorname*{arg\,min}_{\mathbf{v}^{*}\in\mathrm{span}(\mathbf{u}_{1},\cdots,\mathbf{u}_{S})}\left(-\mathbf{v}^{T}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\bm{\theta}}(\bm{\theta})+\frac{1}{2}\mathbf{v}^{T}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\bm{\theta}^{2}}(\bm{\theta})\mathbf{v}\right).$ (30)
@@
$\mathbf{v}^{*}:=\mathbf{U}\mathbf{I}_{P:S}\bar{\mathbf{H}}^{-1}\bar{\mathbf{g}}=\mathbf{U}\mathbf{I}_{P:S}\bm{\eta}^{*},$ (31)
where $\mathbf{U}=\mathrm{Diag}(\mathbf{u})$. So, we recover the method described in Section 4.1.
Appendix C Link with Cauchy’s steepest descent and Newton’s method

### v4

@@ 7 Discussion
We have shown that it is possible to obtain 2nd and 3rd order information about the loss, and that this information can be used to construct per-layer learning rates and an optimization method with interesting properties. However, this optimization method can only be seen as a proof of concept, showing that higher-order derivatives are accessible and can be used to train neural networks, and not as a generic optimizer with excellent results on a wide range of tasks and models. Therefore, we propose future research directions.
@@
Method 1 does not come with a precise convergence rate. The rate proposed in Appendix 0.F (Theorem 0.F.1) gives only a heuristic. Given the convergence rates of Newton’s method and Cauchy’s steepest descent, we can expect to find some in-between convergence rates. Since Cauchy’s steepest method is vulnerable to a highly anisotropic Hessian, it would be valuable to know how much this weakness is overcome by our method.
Accounting for the noise during training.
Our optimization method remains subject to instabilities during training, which is expected for a second-order method, but not acceptable for the end user. In fact, it is very likely that our algorithm would achieve better performance if it were designed from the beginning to work in a stochastic context. Currently, it is designed as if training was done in full batch.
{credits}
7.0.1 Acknowledgements
The project leading to this work has received funding from the French National Research Agency (ANR-21-JSTM-0001 and ANR-19-CHIA-0021). This work was granted access to the HPC resources of IDRIS under the allocation 2024-AD011013762R2 made by GENCI. We thank Julyan Arbel, Michael N. Arbel, Gilles Blanchard and Christophe Giraud for their support.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.
Appendix 0.A Extensions of Pearlmutter’s trick
In this appendix, we show how to use the trick of [28] to compute the terms of the Taylor expansion of $\mathcal{L}$ and the tensor $\mathbf{D}^{d}_{\boldsymbol{\theta}}(\mathbf{u})$ defined in Eqn. (6).
@@ Appendix 0.A Extensions of Pearlmutter’s trick
$\mathcal{L}(\boldsymbol{\theta}+\mathbf{u})=\mathcal{L}(\boldsymbol{\theta})+\sum_{d=1}^{D}\frac{1}{d!}\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})[\mathbf{u},\cdots,\mathbf{u}]+o(\|\mathbf{u}\|^{D}).$ (11)
@@
$\tilde{\mathbf{D}}_{\boldsymbol{\theta}}^{d}(\mathbf{u}):=\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})[\mathbf{u},\cdots,\mathbf{u}]\in\mathbb{R}.$ (12)
@@
$\tilde{\mathbf{D}}_{\boldsymbol{\theta}}^{d-1}(\mathbf{u})=\frac{\mathrm{d}^{d-1}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d-1}}(\boldsymbol{\theta})[\mathbf{u},\cdots,\mathbf{u}]\in\mathbb{R}.$ (13)
@@
$\tilde{\mathbf{D}}_{\boldsymbol{\theta}}^{d}(\mathbf{u})=\left(\frac{\mathrm{d}\tilde{\mathbf{D}}_{\boldsymbol{\theta}}^{d-1}(\mathbf{u})}{\mathrm{d}\boldsymbol{\theta}}\right)^{T}\mathbf{u}.$ (14)
Therefore, at each step $d$, we only have to compute the gradient of a scalar $\tilde{\mathbf{D}}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ according to $\boldsymbol{\theta}\in\mathbb{R}^{P}$, and compute a dot product in the space $\mathbb{R}^{P}$. So, computing $\tilde{\mathbf{D}}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ has a complexity proportional to $d\times P$, and does not require the computation of the full tensor $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}(\boldsymbol{\theta})\in\mathbb{R}^{P^{d}}$.
Fast computation of $\mathbf{D}^{d}_{\boldsymbol{\theta}}(\mathbf{u})$.
We assume that the parameter $\boldsymbol{\theta}$ is represented by a sequence of vectors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$: each coordinate $\theta_{i}$ belongs to exactly one of the $\mathbf{T}^{s}$. In the same way, given a direction $\mathbf{u}\in\mathbb{R}^{P}$ in the space of the parameters, $\mathbf{u}$ can be represented by a sequence of vectors $(\mathbf{U}^{1},\cdots,\mathbf{U}^{S})$.
We want to compute the tensor $\mathbf{D}^{d}_{\boldsymbol{\theta}}(\mathbf{u})\in\mathbb{R}^{S^{d}}$, whose coefficients are:
$(\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u}))_{s_{1},\cdots,s_{d}}$ $=\frac{\partial^{d}\mathcal{L}}{\partial\mathbf{T}^{s_{1}}\cdots\partial\mathbf{T}^{s_{d}}}(\boldsymbol{\theta})[\mathbf{U}^{s_{1}},\cdots,\mathbf{U}^{s_{d}}],$ (15)
@@
Let us assume that $\mathbf{D}^{d-1}_{\boldsymbol{\theta}}(\mathbf{u})$ is available. We can compute the coefficients of $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ as follows:
@@
$(\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u}))_{s_{1},\cdots,s_{d}}=\left(\frac{\partial(\mathbf{D}^{d-1}_{\boldsymbol{\theta}}(\mathbf{u}))_{s_{1},\cdots,s_{d-1}}}{\partial\mathbf{T}^{s_{d}}}\right)^{T}\mathbf{U}^{s_{d}}$ (16)
@@
That way, the tensor $\mathbf{D}_{\boldsymbol{\theta}}^{d}(\mathbf{u})$ can be computed without having to compute the full derivative $\frac{\mathrm{d}^{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{d}}$. In fact, we do not need to store objects of size greater than $S^{d-1}\times P$: the last operation requires storing $\mathbf{D}^{d-1}_{\boldsymbol{\theta}}(\mathbf{u})$, which is of size $S^{d-1}$, and the gradient of each of its elements with respect to $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$, which is of size $P$.
Appendix 0.B Derivation of the optimal learning rates
We consider an update of $\boldsymbol{\theta}$ with one learning rate $\eta_{s}$ for each subset $\mathcal{I}_{s}$ of parameters. Let $\mathbf{I}_{S:P}\in\mathbb{R}^{S\times P}$ be the partition matrix, verifying $(\mathbf{I}_{S:P})_{sp}=1$ if $p\in\mathcal{I}_{s}$ and 00 otherwise, and $\mathbf{I}_{P:S}:=\mathbf{I}_{S:P}^{T}$. We consider an update based on a given direction $\mathbf{u}_{t}$ and we define $\mathbf{U}_{t}:=\mathrm{Diag}(\mathbf{u}_{t})$:
$\boldsymbol{\theta}_{t+1}=\boldsymbol{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta},$ (17)
where $\boldsymbol{\eta}=(\eta_{1},\cdots,\eta_{S})\in\mathbb{R}^{S}$.
@@ Appendix 0.B Derivation of the optimal learning rates
$\mathcal{L}(\boldsymbol{\theta}_{t+1})$ $=\mathcal{L}(\boldsymbol{\theta}_{t}-\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta})$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\boldsymbol{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta}_{t})$
$\hphantom{=}\;+\frac{1}{2}\boldsymbol{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{2}}(\boldsymbol{\theta}_{t})\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}+o(\|\boldsymbol{\eta}\|^{2})$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\boldsymbol{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{g}_{t}$
$\hphantom{=}\;+\frac{1}{2}\boldsymbol{\eta}^{T}\mathbf{I}_{S:P}\mathbf{U}_{t}\mathbf{H}_{t}\mathbf{U}_{t}\mathbf{I}_{P:S}\boldsymbol{\eta}+o(\|\boldsymbol{\eta}\|^{2})$
$=\mathcal{L}(\boldsymbol{\theta}_{t})-\boldsymbol{\eta}^{T}\bar{\mathbf{g}}_{t}+\frac{1}{2}\boldsymbol{\eta}^{T}\bar{\mathbf{H}}_{t}\boldsymbol{\eta}+o(\|\boldsymbol{\eta}\|^{2}),$
@@
Now, we omit the $o(\|\boldsymbol{\eta}\|^{2})$ term and we want to minimize according to $\boldsymbol{\eta}$ the variation of the loss:
$\mathcal{L}(\boldsymbol{\theta}_{t+1})-\mathcal{L}(\boldsymbol{\theta}_{t})\approx\boldsymbol{\Delta}_{2}(\boldsymbol{\eta}):=\mathcal{L}(\boldsymbol{\theta}_{t})-\boldsymbol{\eta}^{T}\bar{\mathbf{g}}_{t}+\frac{1}{2}\boldsymbol{\eta}^{T}\bar{\mathbf{H}}_{t}\boldsymbol{\eta}.$ (19)
We have: $\frac{\mathrm{d}\boldsymbol{\Delta}_{2}}{\mathrm{d}\boldsymbol{\eta}}=-\bar{\mathbf{g}}_{t}+\bar{\mathbf{H}}_{t}\boldsymbol{\eta}$, which is zero if, and only if: $\bar{\mathbf{g}}_{t}=\bar{\mathbf{H}}_{t}\boldsymbol{\eta}$. If this linear system can be inverted, one can choose:
$\boldsymbol{\eta}=\boldsymbol{\eta}_{t}^{*}:=\bar{\mathbf{H}}_{t}^{-1}\bar{\mathbf{g}}_{t}.$ (20)
@@
This method can also be derived by taking the point of view of multivariate optimization. Within our setup, $\boldsymbol{\theta}$ is considered as a tuple of tensors $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. Thus, we want to minimize the loss $\mathcal{L}$ with respect to $(\mathbf{T}^{1},\cdots,\mathbf{T}^{S})$. By abuse of notation, we will write:
$\mathcal{L}(\boldsymbol{\theta})=\mathcal{L}(\mathbf{T}^{1},\cdots,\mathbf{T}^{S}).$ (21)
@@
$\approx-\boldsymbol{\eta}^{T}\bar{\mathbf{g}}+\frac{1}{2}\boldsymbol{\eta}^{T}\bar{\mathbf{H}}\boldsymbol{\eta},$ (25)
where $\boldsymbol{\eta}=(\eta_{1},\cdots,\eta_{S})\in\mathbb{R}^{S}$, $\bar{\mathbf{g}}\in\mathbb{R}^{S}$ is the gradient of $f$ and $\bar{\mathbf{H}}\in\mathbb{R}^{S\times S}$ is the Hessian of $f$:
@@
Finally, one can minimize the order-2 approximation of $f$ (Eqn. (25)) with respect to $\boldsymbol{\eta}$, with any numerical or analytical technique. If done analytically, we roll back to Eqn. (20).
@@
$g(\mathbf{v})=\mathcal{L}(\boldsymbol{\theta}-\mathbf{v})-\mathcal{L}(\boldsymbol{\theta}),$ (28)
@@
$g(\mathbf{v})\approx\tilde{g}(\mathbf{v}):=-\mathbf{v}^{T}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta})+\frac{1}{2}\mathbf{v}^{T}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{2}}(\boldsymbol{\theta})\mathbf{v}.$
@@
$\mathbf{v}^{*}=\operatorname*{arg\,min}_{\mathbf{v}^{*}\in\mathrm{span}(\mathbf{u}_{1},\cdots,\mathbf{u}_{S})}\left(-\mathbf{v}^{T}\frac{\mathrm{d}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}}(\boldsymbol{\theta})+\frac{1}{2}\mathbf{v}^{T}\frac{\mathrm{d}^{2}\mathcal{L}}{\mathrm{d}\boldsymbol{\theta}^{2}}(\boldsymbol{\theta})\mathbf{v}\right).$ (29)
@@
$\mathbf{v}^{*}:=\mathbf{U}\mathbf{I}_{P:S}\bar{\mathbf{H}}^{-1}\bar{\mathbf{g}}=\mathbf{U}\mathbf{I}_{P:S}\boldsymbol{\eta}^{*},$ (30)
where $\mathbf{U}=\mathrm{Diag}(\mathbf{u})$. So, we recover the direction computed in Section 4.
Appendix 0.C Link with Cauchy’s steepest descent and Newton’s method